OpenStudy (anonymous):

Please help, How do I solve this? a) (x^2 -4x)*(x^2 +2x-3)=0 and b)2x^3+7x^2-4x=0

OpenStudy (solomonzelman):

for your first question, you know that either (x^2 -4x) or (x^2 +2x-3) would be equal to zero so solve each (x^2 -4x)=0 AND (x^2 +2x-3) = 0

OpenStudy (solomonzelman):

for your second question, you can factor out of x 2x^3+7x^2-4x=0 x(2x^2+7x-4)=0 divide both sides by x (2x^2+7x-4)=0 2x^2+7x-4=0 can you solve this quadratic ?

OpenStudy (anonymous):

Hey, @SolomonZelman ! Didn't even notice you replied! Um, Can you show me how you solve one of the equation on a) ? And in b), don't i use the ax^2+bx+c=0 and then use -b+/- sqrt(b-4ac)/2ac ?

OpenStudy (solomonzelman):

for a) (x^2 -4x)=0 AND (x^2 +2x-3) = 0 x^2-4x=0 x^2+2x-3=0 x(x)-4(x)=0 (x-1)(x+3)=0 (x-4)x=0 x-4=0 x-1=0 and x+3=0 x=4 x=1 x=-3 for b) 2x^3+7x^2-4x=0 x(2x^2+7x-4)=0 2x^2+7x-4=0 (2x-1)(x+4)=0 2x-1=0 and x+4=0 2x=1 x= 1/2 x= -4

OpenStudy (solomonzelman):

factoring is the best !

OpenStudy (anonymous):

@SolomonZelman Thank you, thank you, thank you!! :D

OpenStudy (solomonzelman):

You welcome !